OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 4 degrees Celsius is.

OpenStudy (anonymous):

C = 5/9 F - 160/9 C + 160/9 = 5/9 F F = 9/5 C + 32

OpenStudy (anonymous):

part 2: F = 9/5 (4) + 32 F = 36/5 + 32 F = 39.2 degrees F.

OpenStudy (diyadiya):

F=9/5*4+32 =36/5+32 =196/5 =39.2 F